OpenStudy (anonymous):

Which of the following are solutions of the quadratic equation below? Check all that apply. 2x^2-2x-2=x A.-1/2 B. 1 C. 3 D. 2 E. -3 F. -4

OpenStudy (anonymous):

subtract from both sides x: 2x2-2x-x-2=0 --------->2x2-3x-2=0 use the quadratic fromula

OpenStudy (fibonaccichick666):

what can you use when you have a quadratic eq?

OpenStudy (anonymous):

\[(3+-\sqrt(9+16))/4\]